A field 100 m by 75 m has a pathway 3 m wide running in thw center horizontally and vertically . please find the cost of paving the pathway at the rate of seven rupees per square foot and planting grass in the rest of the field at the rate of 2 rupees per sq foot.
Answer by Theo(13342)

the pathway across the length of the field measures 3 meters by 100 meters.
the pathway across the width of the field measures 3 meters by 75 meters.
the square in the middle that is contained in both pathways is being counted twice, so it has to be subtracted from the total square feet of the pathway that is being paved.
the total square feet of the pathway is therefore 100 * 3 plus 75 * 3 minus 3 * 3 = 300 plus 225 - 9 = 516 square feet.
the cost to pave this is 7 * 516 = 3612 rupees.
the area of the field that has to have grass planted is equal to 100 * 75 - 516 = 7500 - 516 = 6984 square feet.
the cost to plant grass is 2 per square foot * 6984 = 13968 rupees.
the total cost is 13968 + 3612 = 17580.